Find the complete integral of the p.d.e P2x+Q2x= Pz using Jacobi's method


1
Expert's answer
2021-03-11T13:47:49-0500

Step 1: (Converting the given PDE into the form "f(x, y, z, u_x, u_y, u_z)" = 0.


Set "p = −\dfrac{ux}{uz} and \text q = −\dfrac{u_y}{u_z}" in the given PDE to obtain


"\dfrac{(u_x)^2}{(u_z)^2} x + \dfrac{u_y^2}{ u_z^2} y = z =⇒ xu_x^2 + yu_y^2 − zu_z^2 = 0" .

Thus,


"f(x, y, z, u_x, u_y, u_z) = xu_x^2 + yu_y^2 \u2212 zu_z^2 = 0" .


Step 2: Solving above PDE by Jacobi’s method


Compute "f_{u_x} , f_{u_y} , f_{u_z} , f_x, f_y, f_z"


"f_{u_x} = 2xu_x, f_{u_y} = 2yu_y, f_{u_z} = \u22122zu_z, f_x = u_x^2 , f_y = u_y^2 , f_z = \u2212u_z^2" .

.

Step 2(b): Writing auxiliary equation and solving for "u_x, u_y and u_z" .


The auxiliary equations are given by

"\dfrac{dx}{ f_{u_x}} = \dfrac{dy}{ f_{u_y}} = \dfrac{dz}{ f_{u_z}} = \dfrac{du_x}{ −f_x} = \dfrac{du_y}{ −f_y} = \dfrac{du_z}{ −f_z}"


"\dfrac{dx}{ 2xu_x} = \dfrac{dy}{ 2yu_y} = \dfrac{dz}{ −2zu_z} = \dfrac{du_x}{ −u_x^2} = \dfrac{du_y}{ −u_y^2} = \dfrac{du_z}{ u_z^2}"



Now, "\dfrac{dx }{2xu_x} = \dfrac{du_x}{ −u_x^2} ⇒ \dfrac{u_xdx}{ 2xu_x^2} = −\dfrac{2xdu_x}{ 2xu_x^2}"


"⇒ u_xdx = −2xdu_x \\=

⇒ \dfrac{dx}{ x} = −2 \dfrac{du_x} {u_x} \\

⇒ log x = −2 log(ux) + log(a) \\

⇒ log x + log(u_x^2 ) = log(a)\\ 

⇒ xu_x^2 = a \\

⇒ u_x = (\dfrac{a}{ x} )^\frac{1}{2}"


. Similarly, we get


."yu_y^2 = b ⇒ u_y = (\dfrac{ b}{ y} )^\frac{1}{2}"

and


"u_z = [ \dfrac{(a + b) }{2} ]^\frac{1}{2}"


Step 2(c): Solving the equation ."du = u_xdx + u_ydy + u_zdz"


"du = (\dfrac{a }{x} )^\frac{1}{2} dx + (\dfrac{ b}{ y} )^\frac{1}{2} dy + (\dfrac{ a + b}{ z} )^\frac{1}{2} dz\\

⇒ u = 2(a_x)^ \frac{1}{2} + 2(b_y)^ \frac{1}{2} + 2((a + b)z)^ \frac{1}{2} + c."